Suppose the risk-free rate is 3.56% and an analyst assumes a market risk premium of 5.71%. Firm A just paid a dividend of $1.18 per share. The analyst estimates the β of Firm A to be 1.27 and estimates the dividend growth rate to be 4.15% forever. Firm A has 254.00 million shares outstanding. Firm B just paid a dividend of $1.63 per share. The analyst estimates the β of Firm B to be 0.77 and believes that dividends will grow at 2.01% forever. Firm B has 194.00 million shares outstanding. What is the value of Firm A?
Answer #1

For Firm A,

As per CAPM Model,

Required Rate = Rf + beta(Rm - Rf)

Required Rate = 0.0356 + 1.27(0.0571)

Required Rate = 10.81%

As per Dividend Growth Model,

Stock Price = D0(1 + g)/(r - g)

Stock Price = 1.18(1.0415)/(0.1081 - 0.0415)

Stock Price = $18.45

Value of Firm = Stock price * number of shares outstanding

Value of Firm A = 18.45 * 254

Value of Firm A = $4,686.30 million